OpenStudy (anonymous):

How many different arrangements of the letters in the word "ALGEBRA" are there? 2,520 25 1,260 3.5

OpenStudy (shinalcantara):

since it deals with permutation you'll have the formula when some things are identical as: \[_nP_{n i} = \frac{ n! }{ i! i! ... }\]

OpenStudy (anonymous):

huh

OpenStudy (shinalcantara):

\[_7P_{7(2)} = \frac{ 7! }{ 2! }\]

OpenStudy (shinalcantara):

remember that in the word ALGEBRA you have 2 A's and you have 7 letters all

OpenStudy (shinalcantara):

the 2 A's are the identical things/letters you have. n is the total number of things/letters

OpenStudy (shinalcantara):

you can rewrite it as \[_7P {7(2)} = \frac{ 7! }{ 2! }\] \[_7P {7(2)} = \frac{ 7\times6\times5\times4\times3\times2\times1 }{ 2\times1 }\]